Motional EMF A single rectangular loop is pulled out of a region of uniform magnetic field at a con- stant speed v. Suppose that the loop has a mass m and a resistance R. Its cross- sectional width is L, as shown. 1. Find the induced EMF in the loop. Express your FINAL answer in terms of B, L, and v only 2. Find the induced current, including direction. Express your FINAL answer in terms of B L, v, and R only. 3. Find the electric power dissipated in the loop. Express your FINAL answer in terms of B, L, v, and R only.
Answer #1

1. induced emd = d(flux)/dt

= d (B L x) /dt

= B L dx/dt

= B L v


2. I = e / R = B L v / R counterclockwise


3. P= V I

P = B^2 L^2 v^2 / R